# Depth Matters: Multimodal RGB-D Perception for Robust Autonomous Agents

### v1

@@ IV-C Recurrent Models
The feature extractor models are further differentiated from each other based on 4 RNN backbone choices: LSTM, LTC, CfC and LRC.
The LTC and CfC backbones use a NCP of 19 neurons, i.e., 12 inter-neurons, 6 command neurons, and 1 motor neuron, as LTCs were configured in [1]. We adapted CfCs to this setup as they are the closed-form solutions of LTCs. The LSTM and LRC architectures are configured to have a standard size of 64 features in the hidden state. All RNN’s use a sequence length of 16 frames.
@@ IV-D Training and Metrics
All run configurations include modifiable arguments for the RNN model, input mode, the type of per-channel normalization in the normalization layer, padding, and learning rates. We experimented with Min-Max normalization, Z-Score normalization, and Min-Max followed by Z-Score. During training, we set the maximum epoch number to 100 epochs with a batch size of 20, constrained by the available computational resources. To prevent overfitting, training stops early if the validation loss does not improve for three consecutive epochs. We used Tensorboard to log and analyze all models’ training, validation, and testing results.
We used the Adam optimization algorithm  [28] with learning rates of $\{10^{-3},10^{-4},10^{-5}\}$ and $\epsilon=10^{-7}$ for numerical stability. To predict continuous steering commands from multi-dimensional input, we used the Mean Squared Error (MSE) as loss function. Since MSE produces strictly positive values, losses closest to 0 indicate the most accurate estimations.
@@ V RESULTS
In this section, we summarize our findings on the various model architectures presented in Section IV. First, we describe the results from the passive test using the collected dataset, followed by the deployment results of the best models.
@@ V-A Open-loop Results
TABLE I: Test results, averaged over 3 seeds. Models were trained, validated, and tested on the collected dataset. Loss values are scaled by ×$\times 10^{-3}$. The two best-performing recurrent models for each feature extractor are highlighted.
@@ V-B Closed-loop Results
After training and evaluating the models on the passive dataset, we were interested in how they would adapt to a closed-loop setting when deployed on hardware. Since we had a total of 20 models (5 feature extractors × 4 RNN options), we chose to test only the most promising ones that demonstrated good performance on the dataset.
To ensure a balanced evaluation of both the feature extractors and the RNNs, we created test groups based on both components. Specifically, we selected the best two recurrent models for each feature extractor and the best two feature extractors for each recurrent model across all seeds. The results are presented in Table II.
TABLE II: The summary of the closed-loop testing is presented, with results grouped by different feature extractors on the left and by recurrent models on the right. For each feature extractor, we deployed the two best-performing RNN options. Similarly, for each RNN, we deployed the two best-performing feature extractor options based on the test results from the recorded passive dataset. S denotes success of 5 consecutive laps on the test track.
Feat.-Extr. RNN S
EARLY LSTM
LRC
LATE LSTM
LRC
ZACN LRC
LSTM
DCN LSTM
LRC
RGB LSTM
LRC
RNN Feat.-Extr. S
LSTM LATE
EARLY
LTC LATE
ZACN
CfC LATE
ZACN
LRC LATE
ZACN
Figure 5: Extraction of the map used during closed-loop active testing. Problematic turns are marked with an ’X’ to indicate crashes. Labels specify which model crashed at each point.
We found that this closed-loop setup is very challenging for many of the trained models, with only a subset successfully completing 5 continuous laps without crashing. We marked the sections in Figure 5 where crashes occurred. Interestingly, the LATE feature extractor only succeeded with the CfC model, even though it was not among the two best recurrent models for the LATE type in the passive open-loop test (as shown in Table I). However, we evaluated it because CfCs demonstrated the best performance with this feature extractor method, which highlights the effectiveness of our double grouping approach for feature extractors and recurrent models, respectively. This also illustrates the sim-to-real gap, highlighting the performance mismatch between the open-loop setup and real hardware deployment.
@@
We explored the underlying attention mechanism of these models to gain insight into which parts of the input they focus on the most. To calculate this, we used the VisualBackProp method [29], where bright regions indicate where the model focused its attention. The attention maps are displayed in Figure 6. We found that the EARLYLSTM maintains its focus on the more distant regions, while the other models focus more on the road directly ahead. The attention of the LATECfC, ZACNLRC, and DCNLSTM models tends to explore the right side of the road more, while the RGBLSTM model keeps its attention more evenly in front, without focusing on the turn. We hypothesize that this lack of focus on the turn contributed to the RGBLSTM’s crash.
@@ V-C How Robust Are the Deployed Models?
We selected models that completed 5 consecutive laps on the track without crashing for further evaluation under noisy conditions. These models include EARLYLSTM, LATECfC, ZACNLRC, and DCNLSTM. To test their robustness, we added additional Gaussian noise with a mean of 0 and a variance of 0.1 to the input stream. Only the EARLYLSTM model handled successfully this challenging condition, while the others failed, resulting in crashes at Points 2, 7, and 3, as shown in Figure 5, respectively.
Figure 7: The Structural Similarity Index (SSIM) values for the EARLYLSTM, LATECfC, ZACNLRC, and DCNLSTM models measure the networks’ attention robustness under noise. EARLYLSTM performed the best, with values close to 1, indicating stronger resilience against noisy input than the rest of the models analysed.
@@
Figure 8: Attention maps during closed-loop testing, for input with no-noise and increased Gaussian noise with variance of $\sigma=0.1$, $0.2$ and $0.3$, respectively. Evaluated on the EARLYLSTM model, as it was the best-performing model being able to receive additional increasing input noise. It demonstrates that the model could consistently maintain its attention on the same region.
Figure 9: Robustness against noise during closed-loop testing, evaluated on the EARLYLSTM model, which managed to maintain 100% autonomy without crashing, under increased noise. Here, we illustrate the SSIM index for attention maps between no-noise and increased Gaussian noise with variance of $\sigma=0.1$, $0.2$ and $0.3$, respectively.
Figure 10: The trajectory of the human driver on the test track is shown on the left, while the trajectory of our best model, EARLYLSTM, is shown on the right, both plotted using object tracking. While the model exhibits a smooth trajectory, it does not fully match the driving style of the human driver.
To analyze this quantitatively, we want to see how much the attention of these models gets distorted when extra noise is introduced. For this, we used the Structural Similarity Index (SSIM) [30], which assigns a score between 0 and 1 to measure the similarity between two images. In our case, these images represent pairwise comparisons of attention maps under noise-free and noisy conditions. SSIM values close to 1 indicate that the network’s attention remained unchanged, demonstrating its resilience. The SSIM values between the no-noise and $\sigma=0.1$-variance noise for these models are shown in Figure 7. The significantly higher SSIM value of the EARLYLSTM supports our finding that it can tackle noise with more robust attention, while the other models failed under this condition.
@@
To further assess the robustness of the promising EARLYLSTM model, we gradually increased the noise level to 0.2 and then to 0.3. Impressively, the model remained robust even at these high noise levels. We were particularly interested in how this resilience relates to the model’s attention maps - specifically, how well it maintains focus on key input features without being distracted by noise during steering. This is shown in Figure 8.
@@
The corresponding SSIM values are presented in Figure 9, which demonstrates that even with a high level of noise, the EARLYLSTM managed to keep its attention similarity close to 1, with $86\%,75\%$ and $67\%$, respectively. Although there is a decrease in SSIM values as noise increases, the rate of decrease is not linear, it slows down progressively, with values decreasing by $-14\%$, $-11\%$, and $-8\%$, respectively. This suggests that while the model is affected by noise, it retains much of its attention stability even at higher noise levels. This resilience could be a key factor in its consistent performance under these challenging, noisy conditions.
V-D Comparison with Human Expert Driver
@@
Figure 11: Histogram of steering differences between consecutive predictions, showing the smoothness of steering behavior. Values closer to zero indicate smoother steering. The plots compare the distribution between model predictions and human-commanded steering during closed-loop testing. On the left, the histogram represents real-time inference where $\approx 7$fps were being dropped due to hardware limitations. On the right, we plot the post-hoc inference results. The similarity between the two distributions suggests that frame drops did not significantly affect the model’s overall behavior.
@@
After successfully deploying and testing the EARLYLSTM model under various noisy conditions, we raised the question of how this model compares to human expert driver behavior. To explore this, we recorded a driving session of 5 continuous laps with the human driver on the test track.
To compare the driving trajectory of the human driver and the EARLYLSTM model, we implemented an object tracking algorithm using OpenCV [31] that tracks the agent over time and plots its trajectory on the static test layout. This is shown in Figure 10. Although the model produces a very smooth trajectory, it does not fully replicate the exact human driver’s driving style. We hypothesize that this difference may be due to the experience the human driver possesses. Even though we used 5 different track layouts for training, we believe that incorporating even more layouts into the training pipeline could improve generalization, enabling the model to more closely match the performance of the human driver.
V-E Impact of Frame Drops on Model Performance
@@
During real-time inference, our model experienced frame drops due to hardware limitations, resulting in a drop of approximately 7fps. Despite this, the model successfully completed 5 fully autonomous laps, including three test scenarios with added noise, all in a consistent manner.
@@
To assess whether these frame drops influenced the model’s behavior, we reprocessed the recorded data offline on a virtual machine, ensuring that all frames were used for inference. Figure 11 compares the histogram of pairwise absolute steering differences between the model and the human driver under both conditions: (1) real-time inference with dropped frames and (2) post-hoc inference using all recorded frames. The similarity between the two distributions suggests that even frame drops did not introduce significant erratic behavior or deviation in the model’s predictions.
@@ VI CONCLUSION
Overcoming the sim-to-real gap for autonomous agents poses many challenges. In our paper, we provided an in-depth analysis of our successful deployment of a multimodal RGB-D recurrent controller. During deployment, we found that depth information is crucial – the unimodal RGB feature extractor was insufficient for autonomous navigation, whereas we could successfully deploy a controller for each multimodal RGB-D feature extractor. Our early fusion LSTM agent maintained 100% autonomy in previously unseen scenarios and high resilience to both unpredictable data loss and increased input noise. In order to interpret the model behavior, we employed various numerical and visual techniques, which reinforced our observations from the closed-loop tests.
ACKNOWLEDGMENT
M.F. and R.G. have received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034277. We thank Mihai-Teodor Stănușoiu for his assistance in the data collection as our human driver.

### v3

@@ IV-C Recurrent Models
The models are further differentiated from each other based on 4 RNN backbone choices: LSTM, LTC, CfC and LRC. The LTC and CfC backbones use a NCP of 19 neurons, i.e., 12 inter-neurons, 6 command neurons, and 1 motor neuron, as LTCs were configured in [Lechner_Hasani_Amini_Henzinger_Rus_Grosu_2020]. We adapted CfCs to this setup as they are the closed-form solutions of LTCs. The LSTM and LRC architectures are configured to have a standard size of 64 features in the hidden state. All RNN’s use a sequence length of 16 frames. At 30 fps, 16 frames correspond to $\approx$0.5 seconds of temporal context. This value balances capturing sufficient motion dynamics (turn anticipation) and staying within memory constraints.
@@ IV-D Training and Metrics
All run configurations include arguments for the RNN model, input mode, the type of per-channel normalization in the normalization layer, padding, and learning rates. We experimented with Min-Max normalization, Z-Score normalization, and Min-Max followed by Z-Score. During training, we set the maximum epoch number to 100 epochs with a batch size of 20, constrained by the available computational resources. To prevent overfitting, training stops early if the validation loss does not improve for three consecutive epochs. We used Tensorboard to log and analyze all models’ training, validation, and testing results.
We used the Adam optimization algorithm  [kingma2014adam] with learning rates of $\{10^{-3},10^{-4},10^{-5}\}$ and $\epsilon=10^{-7}$ for numerical stability. To predict continuous steering commands from multi-dimensional input, we used the Mean Squared Error (MSE) as loss function, where values closer to 0 indicate more accurate estimations.
@@ V RESULTS
In this section, we summarize our findings on the various model architectures presented in Section IV. First, we describe the results from the passive test using the collected dataset, followed by deployment results of the best models.
@@ V-A Open-loop Results
TABLE I: Test results, averaged over 3 seeds. Models were trained, validated, and tested on the collected dataset. Loss values are scaled by $\times 10^{-3}$. The two best-performing recurrent models for each feature extractor are highlighted.
@@ V-B Closed-loop Results
After training and evaluating the models on the passive dataset, we were interested in how they would adapt to a closed-loop setting, i.e., when the model’s steering predictions are fed back into the control loop of the vehicle and directly influence subsequent observations. Since we had a total of 20 models (5 feature extractors × 4 RNN options), we chose to first test only the most promising ones that demonstrated good performance on the dataset. To ensure a balanced evaluation of both the feature extractors and the RNNs, we created test groups based on both components. Specifically, we selected the best two recurrent models for each feature extractor and the best two feature extractors for each recurrent model across all seeds.
Figure 5: Extraction of a map used during closed-loop active testing. Problematic turns are marked with an ’X’ to indicate crashes. Labels specify which model crashed at each point.
We first found that a closed-loop setup is very challenging for many models, with only a subset completing 5 continuous laps without crashing on Map 5. No RGB-based models were able to complete this task. We marked the sections in Figure 5 where crashes occurred. Interestingly, the LATE feature extractor only succeeded with the CfC model, even though it was not among the two best recurrent models for the LATE type in the passive open-loop test (as shown in Table I). However, we evaluated it because CfCs demonstrated the best performance with this feature extractor method, which highlights the effectiveness of our double grouping approach for feature extractors and recurrent models, respectively. This also illustrates the sim-to-real gap, highlighting the performance mismatch between the open-loop setup and real hardware deployment.
@@
We explored the underlying attention mechanism of these models to gain insight into which parts of the input they focus on the most. To calculate this, we used the VisualBackProp method [visualbackprop], where bright regions indicate where the model focused its attention. Attention maps from Map 5 are displayed in Figure 6. We found that the EARLYLSTM maintains its focus on the more distant regions, while the other models focus more on the immediate road. The attention of the LATECfC, ZACNLRC, and DCNLSTM models tends to explore the right side of the road more, while the RGBLSTM model keeps its attention evenly in front, without focusing on the turn. We hypothesize that a similar lack of focus on the turn contributed to the RGBLSTM’s crash.
@@ V-C How Robust Are the Deployed Models?
We selected models that completed 5 consecutive laps on the track without crashing for further evaluation under noisy conditions. These models include EARLYLSTM, LATECfC, ZACNLRC, and DCNLSTM. To test their robustness, we added additional Gaussian noise with a mean of 0 and a variance of 0.1 to the input stream. Only the EARLYLSTM model handled successfully this challenging condition, while the others failed, resulting in crashes at Points 2, 7, and 3, as shown in Figure 5, respectively. We hypothesize that this is due to the combination of the smart early fusion mechanism and the optimized LSTM implementation in this resource-constrained setup.
@@
Figure 7: Closed-loop testing with no-noise and increasing Gaussian noise with variance of $\sigma=0.1$, $0.2$ and $0.3$ on the EARLYLSTM model. Top: Attention maps demonstrate that the model could consistently maintain its attention on the same region. Bottom: High SSIM index for attention maps between no-noise and increased Gaussian noise levels.
@@
To further assess the robustness of the promising EARLYLSTM model, we gradually increased the noise level to 0.2 and then to 0.3. Impressively, the model remained robust even at these high noise levels. We were particularly interested in how this resilience relates to the model’s attention maps - specifically, how well it maintains focus on key input features without being distracted by noise during steering. This is shown at the top of Figure 7.
@@
To analyze this quantitatively, we want to see how much the attention of this model gets distorted when extra noise is introduced. For this, we used the Structural Similarity Index (SSIM) [SSIM], which assigns a score between 0 and 1 to measure the similarity between two images. In our case, these images represent pairwise comparisons of attention maps under noise-free and noisy conditions. SSIM values close to 1 indicate that the network’s attention remained unchanged, demonstrating its resilience. This are presented at the bottom of Figure 7, which demonstrates that even with a high level of noise, the EARLYLSTM managed to keep its attention similarity close to 1, with $86\%,75\%$ and $67\%$, respectively. Although there is a decrease in SSIM values as noise increases, the rate of decrease is not linear: it slows down progressively, with values decreasing by $4\%, 11$$1\%$, and $8\%$, respectively. This suggests that while the model is affected by noise, it retains much of its attention stability even at higher noise levels. This resilience could be a key factor in its consistent performance under these challenging, noisy conditions.
V-D Out-of-Distribution Testing
Figure 8: Labeled testing circuit in dark conditions. The figure shows one of the testing tracks with labeled challenges. Different variations were conducted by moving the static obstacles and light sources, repositioning reflective foil on different walls (turns, roundabout), introducing gaps by removing pipe segments or adding/removing the transparent wall (plastic wall).
To understand EARLYLSTM, we constructed a configurable testing circuit with multiple challenging elements (Fig. 8). The environment included reflective surfaces, directed light beams, carpet material, transparent walls, decision-making points: intersection, roundabout, and both static and dynamic obstacles (a second roboracer, controlled by a human). These conditions were not fixed: obstacles and light sources were moved between runs, reflective foil was attached to different wall segments, gaps were opened by removing pipes, and new obstacles were introduced while the vehicle was driving. This ensured that the model was exposed to a wide distribution of untrained situations, testing the ability to generalize in closed-loop deployment.
Figure 9: Comparison of RGBLSTM and EARLYLSTM models in closed-loop obstacle interaction. Overlaid attention maps highlight the difference: while the RGB model remains focused on the blocked route (left), the depth-based EARLYLSTM shifts its attention toward the free path (right).
One of the clearest differences between modalities emerged during obstacle tests. As shown in Fig. 9, EARLYLSTM successfully avoided an unseen obstacle in the roundabout by immediately redirecting its trajectory toward the free branch. In contrast, the RGBLSTM failed to perceive the obstruction, continued to predict a near-straight steering command, and collided with the object. Notably, neither model was trained for obstacle avoidance, yet depth provided an implicit geometric cue that guided safe behavior, while RGB-only perception lacked the necessary information.
Figure 10: Behavior of the EARLYLSTM model at intersections, illustrated with overlaid attention maps. (a) Taking left or right paths depending on its approach. (b) Handling dynamic (other car) and static (white box) obstacles.
Although our training data did not contain intersections or explicit decision-making commands, we observed that the RGB-D models were able to navigate such scenarios in closed loop. Figure 10(a) illustrates two examples where the vehicle approached an intersection from different trajectories. In the first case, the model exited the preceding turn with a shallow angle, allowing visibility of both branches, and ultimately chose the left branch by steering slightly left (0.05 rad). In the second case, the vehicle entered at a sharper angle, giving greater visibility into the right branch, and the model steered accordingly (0.09 rad). This behavior suggests that the controller relied on forward visibility when choosing a path, effectively preferring the branch with more open space ahead. Importantly, this decision-making emerged without explicit supervision for intersections, highlighting the capacity of the RGB-D recurrent models to generalize beyond their training distribution and exploit geometric cues available in the sensory input.
Figure 11: Closed-loop interaction with a dynamic obstacle. Top row: dome-camera view of two roboracer vehicles on the circuit (the front vehicle controlled by a human expert). Bottom row: overlaid attention maps from our EARLYLSTM model with corresponding overlaid attention maps and steering commands.
Introducing a second, human-driven roboracer created dynamic interaction scenarios. Our EARLYLSTM handled these cases without explicit training for multi-agent driving. Figure 11 shows one critical example: as the human-driven vehicle cut in front, the autonomous car adjusted both its attention and steering trajectory in real time. The overlaid attention maps clearly illustrate this adaptation, initially focusing left to overtake, then rapidly shifting right when space became constrained, leading to a smooth avoidance maneuver. This new behavior shows the capacity of our model to exploit RGB-D cues for reactive obstacle avoidance in closed-loop settings.
In Figure 10(b), a static obstacle blocked one branch of the intersection while the human-driven roboracer cuts across the other. In this situation, depth cues indicated no free space to the right, while the human vehicle constrained the left. The model appears to fall back on RGB cues, focusing on the floor texture and up ahead, successfully avoiding both obstacles. This behavior highlights the complementarity of RGB and depth: when depth cues were limited, RGB provided sufficient contextual information to maintain navigation.
Figure 12: Comparison of EARLYLSTM with TinyLidarNet under reflective-surface conditions.
We deployed TinyLidarNet for comparison on our test track. Under normal conditions, TinyLidarNet performed reliably, producing smooth trajectories and avoiding obstacles. However, when reflective foil was placed on a wall in Section 1 (labeled in Figure 8), its predictions collapsed: steering commands became inconsistent, and the vehicle ultimately crashed into the barrier (Figure 12). Quantitatively, TinyLidarNet’s steering smoothness degraded severely: the maximum difference between consecutive steering commands dropped from 0.23 (foil covered) to 0.09 (foil exposed). The low value indicates the network was repeating nearly identical predictions, effectively freezing and failing to adapt to the turn. This aligns with known limitations of LiDAR when encountering specular surfaces, which cause signal loss or spurious returns. In contrast, EARLYLSTM was unaffected by the foil and navigated the same section without difficulty. These results validate our choice of RGB-D sensing as a more robust alternative for our scenario.
V-E Comparison with Human Expert Driver
Figure 13: The trajectory of the human driver on the test track is shown on the left, while the trajectory of our best model, EARLYLSTM, is shown on the right, both plotted using object tracking. While the model exhibits a smooth trajectory, it does not fully copy the driving style of the human driver.
@@ V-E Comparison with Human Expert Driver
Figure 14: Histogram of steering differences between consecutive predictions, illustrating steering smoothness. Values near zero indicate smoother steering. Left: real-time inference ($\approx 7$fps dropped due to hardware limits). Right: post-hoc inference. Similar distributions suggest frame drops had minimal effect on overall behavior.
@@
After successfully deploying and testing the EARLYLSTM model under various noisy conditions, we raised the question of how this model compares to human expert driver behavior. To explore this, we recorded a driving session of 5 continuous laps with the human driver on Map 5.
To compare the driving trajectory of the human driver and the EARLYLSTM model, we implemented an object tracking algorithm using OpenCV [opencv_library] that tracks the agent over time and plots its trajectory on the static test layout (Fig. 13). Although the model produces a smooth trajectory, it does not fully replicate the human driver’s style. We believe this difference is largely due to the limited amount of expert data available for training. With additional demonstrations from the same driver, the model would likely better capture the nuances of their behavior. Nevertheless, even with only a few minutes of driving data per track, the model already demonstrates effective closed-loop decision-making, highlighting its ability to generalize beyond the training distribution.
V-F Impact of Frame Drops on Model Performance
@@
During real-time inference, our model experienced frame drops due to hardware limitations, resulting in a drop of approximately 7fps. Despite this, the model successfully completed 5 fully autonomous laps on Map 5, including three test scenarios with added noise, and multiple configurations and scenarios on Map 8.
@@
To assess whether these frame drops influenced the model’s behavior, we reprocessed the recorded data offline on a virtual machine, ensuring that all frames were used for inference. Figure 14 compares the histogram of pairwise absolute steering differences on Map 5 between the model and the human driver under both conditions: (1) real-time inference with dropped frames and (2) post-hoc inference using all recorded frames. The similarity between the distributions suggests that even frame drops did not introduce significant erratic behavior or deviation in the model’s predictions.
@@ VI CONCLUSION
Overcoming the sim-to-real gap for autonomous agents poses many challenges. In our paper, we provided an in-depth analysis of our successful deployment of a multimodal RGB-D recurrent controller. During deployment, we found that depth information is crucial – the unimodal RGB feature extractor was insufficient for autonomous navigation, whereas we could successfully deploy a controller for each multimodal RGB-D feature extractor. Our early fusion LSTM agent maintained autonomy in previously unseen scenarios and high resilience to data loss and increased input noise. In order to interpret the model behavior, we employed various numerical and visual techniques, which reinforced our observations from the closed-loop tests.